OpenStudy (anonymous):

. If l is the side of a square and dl/dt =0.2 m/s ,then what is the rate of increase in area when l=5m ? i)2m²/s ii) 1m²/s iii)4m²/s iv)3m²/s

hartnn (hartnn):

Area of square A = l^2 Rate of Increase in Area = dA/dt = (dA/dl)* (dl/dt) = ....?

OpenStudy (anonymous):

Yes!!! i GOT AN answer. we need to find ; dA/dt=? Now, dA/dt=d(l^2)/dt =2l*dl/dt = 2*5*0.2 = 2 hence, answer is 1

hartnn (hartnn):

thats correct! good work :)